# Scalable Composite Likelihood Estimation of Probit Models with Crossed Random Effects

### v2

@@ 3.1 Three misspecified probit models
$L_{\text{row}}(\tau^{2}_{A})=\tau_{A}^{-R}\prod_{i=1}^{R}\int_{\mathbb{R}}L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})\varphi\Bigl{(}\frac{u_{i}}{\tau_{A}}\Bigr{)}\,\mathrm{d}u_{i},$ (9)
where $L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})$ is the conditional likelihood of $\hat{\bm{\gamma}}_{A}=\hat{\bm{\gamma}}(1+\tau^{2}_{A})^{1/2}$ given $u_{i},$
$L_{i\text{\tiny$\bullet$}}(\hat{\bm{\gamma}}_{A}\!\mid\!u_{i})=\prod_{j|i}\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{A}+u_{i})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{A}-u_{i})^{1-y_{ij}},$ (10)
where $j|i={\{j:(i,j)\in\mathcal{S}\}}$ is the set of indices $j$ such that $(\bm{x}_{ij},Y_{ij})$ is observed. Notice that the row-wise likelihood is a function of $\tau^{2}_{A}$ only, because we fix $\bm{\gamma}$ at the estimate $\hat{\bm{\gamma}}$ obtained from the maximization of the all likelihood at the previous step. Rows with a single observation do not contribute to estimation of $\tau^{2}_{A}$ because $\Pr(Y_{ij}=1)=\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}),$ which does not depend on $\tau^{2}_{A}.$
Computation of the row-wise likelihood requires one to approximate $R$ univariate integrals of a standard hierarchical probit model. Since $R$ is large in the applications that motivate this work, then accurate approximation of the one-dimensional integrals is crucial. Otherwise, the accumulation of approximation errors could induce serious biases in the estimation of $\tau^{2}_{A}$. Our intensive numerical studies indicated that accurate approximation of the integrals is obtained by the well-established adaptive Gaussian quadrature (Liu and Pierce, 1994) with a suitable choice of the number of quadrature nodes. Following Jin and Andersson (2020) and Stringer and Bilodeau (2022), we set the number of quadrature nodes equal to $k(R)=\lceil(3/2)\log_{2}(R)-2\rceil,$ a number rather larger than is typically considered in the GLMM literature. For example, in the Stitch Fix application discussed in Section 5, we have $R=741,221$ and therefore we have used $k=28$ quadrature nodes.
Reversing rows and columns, $\tau^{2}_{B}$ is estimated from the column-wise likelihood
@@
$L_{\text{col}}(\tau^{2}_{B})=\tau_{B}^{-C}\prod_{j=1}^{C}\int_{\mathbb{R}}L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})\varphi\Bigl{(}\frac{v_{j}}{\tau_{B}}\Bigr{)}\,\mathrm{d}v_{j},$ (11)
@@
where $L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})$ is the conditional likelihood of $\hat{\bm{\gamma}}_{B}=\hat{\bm{\gamma}}(1+\tau^{2}_{B})^{1/2}$ given $v_{j},$
$L_{\text{\tiny$\bullet$}j}(\hat{\bm{\gamma}}_{B}\!\mid\!v_{j})=\prod_{i|j}\Phi(\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{B}+v_{j})^{y_{ij}}\Phi(-\bm{x}_{ij}^{\mathsf{T}}\hat{\bm{\gamma}}_{B}-v_{j})^{1-y_{ij}}$
with $i|j={\{i:(i,j)\in\mathcal{S}\}}.$ Finally, the estimates under the original parameterization are obtained by inverting the equations in (7),
$\hat{\bm{\beta}}=\hat{\bm{\gamma}}(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2},\quad\hat{\sigma}_{A}^{2}=\frac{\hat{\tau}_{A}^{2}(1+\hat{\tau}^{2}_{B})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}\quad\text{and}\quad\hat{\sigma}_{B}^{2}=\frac{\hat{\tau}_{B}^{2}(1+\hat{\tau}^{2}_{A})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}.$
@@
Note that by the definitions in (7), $\tau^{2}_{A}\tau^{2}_{B}<1$. In our computations, we never encountered a setting where $\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}>1$ [should be $\geq$ not $>$] and so our estimates $\hat{\sigma}^{2}_{A}$ and $\hat{\sigma}^{2}_{B}$ were never negative. In case of a violation, we suggest taking $\hat{\sigma}^{2}_{A}=\hat{\sigma}^{2}_{B}=0$ as is commonly done for negative variance component estimates in random effect settings.
The ARC method offers a significant computational saving compared to ordinary maximum likelihood estimation for two reasons. In addition to the simplified integration, the reparameterization used by the ARC method conveniently separates estimation of the parameters: regression parameters are estimated from a standard probit model, whereas estimation of the dependence parameters require two distinct univariate optimizations of the row- and column-wise likelihoods.
@@
In the next Theorem, we prove the weak consistency of the maximizer $\hat{\bm{\gamma}}$ of the ‘all’ likelihood (8) to the value $\bm{\gamma}$ of equation (7). First we introduce some notation. Let $Z_{ij}=1$ if $(\bm{x}_{ij},Y_{ij})$ is observed and take $Z_{ij}=0$ otherwise. The number of observations in row $i$ is $N_{i\text{\tiny$\bullet$}}=\sum_{j=1}^{C}Z_{ij}$ and similarly column $j$ has $N_{\text{\tiny$\bullet$}j}=\sum_{i=1}^{R}Z_{ij}$ observations in it. Let $\epsilon_{R}=\max_{i}N_{i\text{\tiny$\bullet$}}/N$ and $\epsilon_{C}=\max_{j}N_{\text{\tiny$\bullet$}j}/N$. In our motivating settings it is reasonable to assume that $\max(\epsilon_{R},\epsilon_{C})\to 0$ as $N\to\infty$. When we need a double sum over all pairs of data we use index $r$ as a ‘second $i$’ and $s$ as a ‘second $j$’. We take $Z_{ij}$ to be deterministic with at most one observation for any $(i,j)$ pair. In our motivating applications that would either never happen or be very rare. If rare, one could use the most recent observation only. The $\bm{x}_{ij}$ are not dependent on the $Z_{ij}$. Finally, $Y_{ij}$ are sampled from their probit distribution conditionally on $\bm{x}_{ij}$.
Theorem 1.
@@
Let $Y_{ij}\in\{0,1\}$ follow the crossed random effects probit model (1) with true value $\bm{\gamma}_{0}$ for the parameter $\bm{\gamma}=\bm{\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ Let the number of observations $N\to\infty$ while $\max(\epsilon_{R},\epsilon_{C})\to 0$. Let $\bm{x}_{ij}\in\mathbb{R}^{p}$ satisfy
@@
    $N^{-1}\sum_{ij}Z_{ij}\bm{x}_{ij}\bm{x}_{ij}^{\mathsf{T}}\to V\in\mathbb{R}^{p\times p},$ where $V$ is positive definite;
@@
    there is no nonzero vector $\bm{v}\in\mathbb{R}^{p}$ such that $\bm{v}^{\mathsf{T}}\bm{x}_{ij}\geq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=1$ and $\bm{v}^{\mathsf{T}}\bm{x}_{ij}\leq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=0$.
Let $\hat{\bm{\gamma}}\in\mathbb{R}^{p}$ be any maximizer of (8). Then for any $\epsilon>0$
$\Pr(\|\hat{\bm{\gamma}}-\bm{\gamma}_{0}\|>\epsilon)\to 0,\;\text{as $N\to\infty$}.$
The proof is reported in Section S1 of the Supplementary Material and we obtained it by adapting the proof strategy of Lumley and Mayer Hamblett (2003) to our setting.
Now we consider consistent estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the row and column likelihoods, respectively. Those provide consistent estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ with which one can then adjust the consistent estimate of $\bm{\gamma}$ to get a consistent estimate of $\bm{\beta}$.
Theorem 2.
@@
Our proof of this theorem is also reported in Section S1 of the Supplementary Material. The proof uses the subset argument of Jiang (2013) to show Cramer consistency of the MLE row likelihood and equivalently of the column likelihood. If one does not find Cramer consistency sufficient, then it is possible to construct a consistent estimator using all of the data from a subset of rows with a large number of observations. In Section S1 of the Supplementary Material we show how to construct such a consistent estimator and we give conditions under which the number of large rows will diverge to infinity as $N\to\infty$. See Theorem 3 of the Supplementary Material. We prefer our ARC estimator to an approach using just large rows, because it would be awkward to have to decide in practice which rows to use, and because we believe that there is valuable information in the other, smaller, rows.

### v4

@@ 2.1 Preliminaries
$L({\theta})=\sigma_{A}^{-R}\sigma_{B}^{-C}\int_{\mathbb{R}^{R+C}}L({\beta}\mid{a},{b})\prod_{i=1}^{R}\varphi\Bigl{(}\frac{a_{i}}{\sigma_{A}}\Bigr{)}\prod_{j=1}^{C}\varphi\Bigl{(}\frac{b_{j}}{\sigma_{B}}\Bigr{)}\,\mathrm{d}{a}\,\mathrm{d}{b},$ (3)
where $\varphi(\cdot)$ is the standard normal probability density function, and $L({\beta}\mid{a},{b})$ is the conditional likelihood of ${\beta}$ given the random effects. The conditional likelihood we need is
$L({\beta}\mid{a},{b})=\prod_{(i,j)\in\mathcal{S}}\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j})^{y_{ij}}\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}-a_{i}-b_{j})^{1-y_{ij}},$
and $L(\theta)$ is commonly called the marginal likelihood.
The first-order Laplace approximation is a standard approach to approximate the integral in the marginal likelihood. The Laplace algorithm maximizes the logarithm of the integrand in the marginal likelihood (3) over ${a}\in\mathbb{R}^{R}$ and ${b}\in\mathbb{R}^{C}$ for fixed ${\theta}$. It then multiplies the maximum value of the integrand by $\det\{H^{-1/2}(\theta)\}$ where $H(\theta)\in\mathbb{R}^{(R+C)\times(R+C)}$ is the Hessian of the log integrand with respect to ${a}$ and ${b}$ for fixed ${\theta}$. The result is an approximate marginal likelihood $\tilde{L}({\theta})$ that is optimized to get $\hat{{\theta}}$. See for example Ogden (2021) or Shun & McCullagh (1995). If the square root of the inverse Hessian is computed by standard methods then that alone has cost of $\Omega(N^{3/2})$ by the argument for generalised linear models discussed in §1. Similarly, if the inner optimization is done using Newton steps, that will have a cost that is $\Omega(N^{3/2})$ per iteration. We return to this issue in §3.2 where a Laplace approximation shows superlinear cost that is $o(N^{3/2})$.
Even if the Laplace approximation were computable for large $N$, it does not provide asymptotically valid results in our context because the size of the likelihood integral corresponds to the number of random effects $\Omega(N^{1/2})$ and therefore grows too fast with the sample size to ensure consistent results. See Shun & McCullagh (1995), Ogden (2021) and Tang & Reid (2024) for a discussion of the conditions that must be satisfied to make the Laplace approximation reliable when the size of the integral grows with the sample size.
2.2 Scalable composite likelihood inference
Our approach for a consistent and scalable estimator in high-dimensional probit models with crossed-random effects combines estimates for three misspecified probit models. Each of them is constructed through the omission of some random effects. By combining (1) and (2) we find that marginally
$\mathrm{pr}(Y_{ij}=1)=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}),$ (4)
for ${\gamma}={\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}$. The proposed method begins with estimation of ${\gamma}$ from the naïve model (4) that omits both of the random effects through maximization of the likelihood
$L_{\text{all}}({\gamma})=\prod_{(i,j)\in\mathcal{S}}\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})^{y_{ij}}\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma})^{1-y_{ij}}.$ (5)
Maximization of this likelihood requires neither high dimensional integrals nor expensive algebra and we see it taking $O(N)$ computation in our examples. We then need estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ to get the scale right. Our analysis of (4) will also account for within-row and within-column correlations among the $Y_{ij}$. While $\mathrm{sign}({\gamma}_{k})=\mathrm{sign}(\beta_{k})$ $(k=1,\dots,p),$ confidence intervals for ${\gamma}_{k}$ based on model (4) would be naïve if they did not account for the dependence among the responses.
Consider the reparameterization ${\psi}=({\gamma}^{\mathrm{\scriptscriptstyle T}},\tau^{2}_{A},\tau^{2}_{B})^{\mathrm{\scriptscriptstyle T}},$ where
${\gamma}=\frac{{\beta}}{(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}},\quad\tau^{2}_{A}=\frac{\sigma^{2}_{A}}{1+\sigma^{2}_{B}},\quad\tau^{2}_{B}=\frac{\sigma^{2}_{B}}{1+\sigma^{2}_{A}}.$ (6)
After dividing ${x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\beta}+a_{i}+b_{j}+\varepsilon_{ij}$ by $(1+\sigma^{2}_{A})^{1/2}$ or by $(1+\sigma^{2}_{B})^{1/2}$, model (2) implies that
$\mathrm{pr}(Y_{ij}=1\mid{a})$ $=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{A}+u_{i}),$ (7)
$\mathrm{pr}(Y_{ij}=1\mid{b})$ $=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{B}+v_{j}),$ (8)
where ${u}\sim{N}(0,\tau^{2}_{A}I_{R})$ and ${v}\sim{N}(0,\tau^{2}_{B}I_{C})$ for ${\gamma}_{A}={\gamma}(1+\tau^{2}_{A})^{1/2}$ and ${\gamma}_{B}={\gamma}(1+\tau^{2}_{B})^{1/2}$. Given $\hat{{\gamma}}$, the maximizer of (5), we proceed with estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the two probit models (7) and (8) that each omit one of the random effects. Fitting models (7) and (8) involve simpler integrals than (1) since the latent variable representations of these models
$Y_{ij}=1\{{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{A}+u_{i}+\varepsilon_{ij}>0\},\quad Y_{ij}=1\{{x}_{ij}^{\mathrm{\scriptscriptstyle T}}{\gamma}_{B}+v_{j}+\varepsilon_{ij}>0\},$
have hierarchical (not crossed) error structures. This is where we are able to replace the $(R+C)$-dimensional integral (3) by $R+C$ univariate ones. Model (7) is fitted by maximizing the row-wise likelihood
@@ 2.2 Scalable composite likelihood inference
$L_{\text{row}}(\tau^{2}_{A})=\tau_{A}^{-R}\prod_{i=1}^{R}\int_{\mathbb{R}}L_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}_{A}\mid u_{i})\varphi\Bigl{(}\frac{u_{i}}{\tau_{A}}\Bigr{)}\,\mathrm{d}u_{i},$ (9)
@@
where $L_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}_{A}\mid u_{i})$ is the conditional likelihood of $\hat{{\gamma}}_{A}=\hat{{\gamma}}(1+\tau^{2}_{A})^{1/2}$ given $u_{i},$
$L_{i\text{\tiny$\bullet$}}(\hat{{\gamma}}_{A}\mid u_{i})=\prod_{j\mid i}\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}_{A}+u_{i})^{y_{ij}}\Phi(-{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}_{A}-u_{i})^{1-y_{ij}},$ (10)
where $j\mid i={\{j:(i,j)\in\mathcal{S}\}}$ is the set of indices $j$ such that $({x}_{ij},Y_{ij})$ is observed. The row-wise likelihood (9) is a product of $R$ one-dimensional integrals and is a function of $\tau^{2}_{A}$ only, because we fix ${\gamma}$ at the estimate $\hat{{\gamma}}$ obtained from the maximization of the all likelihood at the previous step. Rows with a single observation do not contribute to estimation of $\tau^{2}_{A}$ because $\mathrm{pr}(Y_{ij}=1)=\Phi({x}_{ij}^{\mathrm{\scriptscriptstyle T}}\hat{{\gamma}}),$ which does not depend on $\tau^{2}_{A}.$ Reversing the roles of the rows and columns, we get a column-wise likelihood $L_{\text{col}}$ which we maximize to get an estimate $\hat{\tau}^{2}_{B}$ of $\tau^{2}_{B}$.
Finally, we invert the equations in (6) to get
$\hat{{\beta}}=\hat{{\gamma}}(1+\hat{\sigma}_{A}^{2}+\hat{\sigma}_{B}^{2})^{1/2},\quad\hat{\sigma}_{A}^{2}=\frac{\hat{\tau}_{A}^{2}(1+\hat{\tau}^{2}_{B})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}},\quad\hat{\sigma}_{B}^{2}=\frac{\hat{\tau}_{B}^{2}(1+\hat{\tau}^{2}_{A})}{1-\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}}.$
@@
Note that by the definitions in (6), $\tau^{2}_{A}\tau^{2}_{B}<1$. In our computations, we never encountered a setting where $\hat{\tau}^{2}_{A}\hat{\tau}^{2}_{B}\geq 1$ and so our estimates $\hat{\sigma}^{2}_{A}$ and $\hat{\sigma}^{2}_{B}$ were never negative.
We call our method ‘all-row-column’. The name comes from model (4) that uses all the data at once, model (7) that combines likelihood contributions from within each row and model (8) that combines likelihood contributions from within each column. Figure 1 illustrates those models for a data set of $N=39$ observations in $R=10$ rows and $C=10$ columns. In each of three misspecified probit models, points in different boxes are assumed to be independent. The left panel of Figure 1 illustrates the ‘all’ model with independent data. The next two illustrate the ‘row’ and ‘column’ hierarchical models. We combine fits from these three misspecified models to get consistent scalable formulas despite the dependencies involved. This approach is a new form of composite likelihood (Lindsay, 1988; Varin et al., 2011). An earlier version of composite likelihood that was applied to crossed random effects is discussed in Bellio & Varin (2005). They considered a standard pairwise likelihood which is however not scalable and thus inappropriate for our problem. Instead, our all-row-column method has $O(N)$ cost per iteration and we get $O(N)$ cost empirically.
Figure 1: The all-row-column method for $N=39$ observations (marked $\bullet$) in $R=10$ rows and $C=10$ columns.
Computation of the row- and column-wise likelihoods require one to approximate up to $R$ and $C$ univariate integrals of a standard hierarchical probit model. Since $R$ and $C$ are large in the applications that motivate this work, then accurate approximation of the one-dimensional integrals is crucial. Otherwise, the accumulation of approximation errors could induce serious biases in the estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$. Our intensive numerical studies indicated that accurate approximation of the integrals is obtained by the well-established adaptive Gauss-Hermite quadrature (e.g., Liu & Pierce, 1994) with a suitable choice of the number of quadrature nodes.
@@
In the next Theorem, we prove the weak consistency of the maximizer $\hat{{\gamma}}$ of the ‘all’ likelihood (5) to the value ${\gamma}$ of equation (6). First we introduce some notation. Let $Z_{ij}=1$ if $({x}_{ij},Y_{ij})$ is observed and take $Z_{ij}=0$ otherwise. The number of observations in row $i$ is $N_{i\text{\tiny$\bullet$}}=\sum_{j=1}^{C}Z_{ij}$ and similarly column $j$ has $N_{\text{\tiny$\bullet$}j}=\sum_{i=1}^{R}Z_{ij}$ observations in it. Let $\epsilon_{R}=\max_{i}N_{i\text{\tiny$\bullet$}}/N$ and $\epsilon_{C}=\max_{j}N_{\text{\tiny$\bullet$}j}/N$. We assume that $\max(\epsilon_{R},\epsilon_{C})\to 0$ as $N\to\infty$. We take $Z_{ij}$ to be deterministic with at most one observation for any $(i,j)$ pair. In our motivating applications, we would seldom if ever have multiple observations for one $(i,j)$ pair. Even then, one might only use the most recent of those observations. The ${x}_{ij}$ are not dependent on the $Z_{ij}$. Finally, $Y_{ij}$ are sampled from their probit distribution conditionally on ${x}_{ij}$.
Theorem 1
@@
Let $Y_{ij}\in\{0,1\}$ follow the crossed random effects probit model (1) with true value ${\gamma}_{0}$ for the parameter ${\gamma}={\beta}/(1+\sigma^{2}_{A}+\sigma^{2}_{B})^{1/2}.$ Let the number of observations $N\to\infty$ while $\max(\epsilon_{R},\epsilon_{C})\to 0$. Let ${x}_{ij}\in\mathbb{R}^{p}$ satisfy
@@
    $N^{-1}\sum_{ij}Z_{ij}{x}_{ij}{x}_{ij}^{\mathrm{\scriptscriptstyle T}}\to V\in\mathbb{R}^{p\times p},$ where $V$ is positive definite;
@@
    there is no nonzero vector ${v}\in\mathbb{R}^{p}$ such that ${v}^{\mathrm{\scriptscriptstyle T}}{x}_{ij}\geq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=1$ and ${v}^{\mathrm{\scriptscriptstyle T}}{x}_{ij}\leq 0$ for all $(i,j)$ with $Z_{ij}=1$ and $y_{ij}=0$.
Let $\hat{{\gamma}}\in\mathbb{R}^{p}$ be any maximizer of (5). Then for any $\epsilon>0$
$\mathrm{pr}(\|\hat{{\gamma}}-{\gamma}_{0}\|>\epsilon)\to 0,\;\text{as $N\to\infty$}.$
The proof is reported in Section S1 of the Supplementary Material and we obtained it by adapting the proof strategy of Lumley & Mayer Hamblett (2003) to our setting. The balance condition that no single row or column have a non-vanishing fraction of data is much weaker than the norm for the crossed random effects literature.
Now we consider consistent estimation of $\tau^{2}_{A}$ and $\tau^{2}_{B}$ from the row and column likelihoods, respectively. Those provide consistent estimates of $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$ with which one can then adjust the consistent estimate of ${\gamma}$ to get a consistent estimate of ${\beta}$.
Theorem 2
@@
Our proof of this theorem is also reported in §S1 of the Supplementary Material. The proof uses the subset argument of Jiang (2013) to show Cramer consistency of the maximum row likelihood estimator of $\tau^{2}_{A}$ and equivalently of the maximum column likelihood estimator of $\tau^{2}_{B}$.